OpenStudy (anonymous):

A spherical balloon is inflated so that its volume is increasing at the rate of 2 ft3/min. How rapidly is the diameter of the balloon increasing when the diameter is 1.7 feet?

OpenStudy (anonymous):

v = 4/3*pi*r^3 dv/dt = 4*pi*r^2*dr/dt when d = 1.7, r = 0.85, and dv/dt = 2: 2 = 4*pi*(0.85)^2*dr/dt thus dr/dt = 1/(2pi*(0.85)^2)

OpenStudy (anonymous):

= 0.2203 ft/min

OpenStudy (anonymous):

.2202836583 was my initial answer too, but apparently wrong. I pretty much did the same thing you did

OpenStudy (anonymous):

ohhh, lol, you're asked for the rate of change of the diameter, not the radius

OpenStudy (anonymous):

wait, nvm. it's 2/(2*pi*((.85)^2))

OpenStudy (anonymous):

hence .440567

OpenStudy (anonymous):

we both read the question incorrectly :)

OpenStudy (anonymous):

yes, you're asked for d/dt(2r) = 2 * dr/dt

OpenStudy (anonymous):

yeah, I blame the distractions...